OpenStudy (anonymous):

Let f(x)=x+5 and g(x)=x^2. Find (g x f)(4) a.9 b.81 c.20 d.1

OpenStudy (anonymous):

(g x f)(4)=g(f(4)=g(4+5)=g(9)=9^2=81

OpenStudy (anonymous):

Thank you!!!

OpenStudy (unklerhaukus):

I think you wrote the question wrong \[(g\times f)(4)=\left((x^2)\times(x+5)\right)\big|^{x=4}=144\] \[(g\circ f)(4)=g(f(x))=\left(x+5\right)^2\big|^{x=4}=9^2=81\]

OpenStudy (anonymous):

welcome